Please please IITians help me in this question. A sphere of radius R is centred at origin and is placed in uniform electrostatic field of intensity E0/√2 i^+E0/√2j^.A quater of the sphere ,lying in the 3rd quadrant is cut out.Find the flux through the remaining curved surface of sphere
Ans-E0πR2/√2

Grade:12th pass

2 Answers

Arun
3 years ago
ϕplain+ϕcurve=0 or ϕplain=ϕcurve
A1=2πR2i^,A2=2πR2j^
E=Ecos45i^+Esin45j^
2Ei^+2Ej^
ϕ=E(A1+A2)
2E2πR22E2πR2=2πR2E
Thus is the flux entering. So flux is 2πR2E
Vikas TU
3 years ago
ϕpla∈+ϕcurve=0
or
ϕpla∈=−ϕcurve
A→1=−πR^2/2iˆ,
A→2=−πR^2/2jˆ
E→=Ecos45∘iˆ+Esin45∘jˆ
=E/√2iˆ+E/√2jˆ
ϕ=E→⋅(A→1+A→2)
=−E/√2πR2^/2−E/√2 πR^2/2=−πR^2E/√2
This is the flux entering. So flux is πR^2E/√2.